OpenStudy (anonymous):

someone please help! Use the table of the function f(x) = x^4 - 2x^3 to answer this question: x f(x) -1 3 0 0 1 -1 2 0 3 27

OpenStudy (sidsiddhartha):

and what are u willing to find out? :P

OpenStudy (mayankdevnani):

@andy11 just plug given values of x in the equation and then match your answer

OpenStudy (mayankdevnani):

let's take an example :- x=-1 and f(x)=3 your equation:- \[\huge \bf f(x)= x^4-2x^3\] plug x=-1, \[\huge \bf f(x)=1-2(-1)=1+2=3\] So,it satisfies.

OpenStudy (mayankdevnani):

similarly,can you do yourself? @andy11

OpenStudy (anonymous):

ok yeah i kinda get it. so for the next one i would try i would replace the -1 with 0?

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

now,you get it

OpenStudy (anonymous):

yes thank you! ill reply back if i did something worng and need help. thanks again

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (anonymous):

ok so i got 1 for the answer. is that correct or did i do somthing wrong

OpenStudy (anonymous):

A plot from Mathematica is attached.

OpenStudy (acxbox22):

all of the x values satisfy the equation so if you plug all the x values in the table into the equation, you get f(x)